how do you figure out this problem?
5x-5y-13=0
x intercept =
y intercept =

Answer by stanbon(75887)

5x-5y-13=0
x intercept =
Let y = 0 ; then x = -13/5
-------------------------------
y intercept =
Let x = 0 ; then y = 13/5
================================
Cheers,
Stan H.